# VinePPO: Unlocking RL Potential For LLM Reasoning Through Refined Credit Assignment

### v1

@@ 8 Discussion
Accurate credit assignment has profound implications on the performance of RL tuning of LLMs. As we’ve demonstrated, standard PPO, despite outperforming most RL-free baselines, suffers from suboptimal value estimation. More importantly, its scaling behavior is concerning; PPO struggles with increasingly diverse trajectories and tends to perform worse as tasks become more complex.
VinePPO, on the other hand, is a viable alternative. As shown in Section 6.4, it offers lowered memory requirements and better performance with the same computational budget. VinePPO could also be a particularly attractive option for frontier LLMs as even doubling the post-training compute is negligible compared to their pre-training costs (Ouyang et al., 2022) [^7] For example, InstructGPT used nearly 60 times more compute for pre-training (Ouyang et al., 2022).. Given the major investments in pre-training compute and data collection of these models, it is imperative for model developers to employ post-training methods that provide more accurate updates, avoiding the high-variance adjustments caused by inferior credit assignment. Additionally, VinePPO offers a straightforward scaling axis: increasing the number of MC samples directly enhances performance with additional compute. Unlike recent approaches that focus on increasing inference-time compute to boost performance (OpenAI, 2024; Bansal et al., 2024), VinePPO’s training compute is amortized over all future inferences. Note that the computational workload of VinePPO is highly parallelizable with linear scalability, making it well-suited for large-scale training.
The unique properties of the language environment are what enabled VinePPO to be viable credit assignment option; it may have limited practical use in traditional deep RL policy gradient methods. This suggests that adapting RL techniques to LLMs requires careful consideration and perhaps a reevaluation of underlying assumptions. Overall, our work highlights the potential of well-tuned RL finetuning strategies with proper credit assignment, and we hope it encourages further research into optimizing RL post-training pipelines for LLMs.
@@ C.4 Evaluation
C.5 Baselines
@@
Therefore, maximizing log-likelihood training on correct responses is equivalent to train with policy gradient without precise credit assignment, such as without advantages for specific actions. In our experiments, we observe the impact of this limitation in both Figure C.3 and Figure C.4 where RestEM overfits on the training data.
@@ C.6 Hyperparameters
In this section, we present a comprehensive overview of the hyperparameters used in our experiments. It’s important to note that the number of training samples was carefully selected to ensure that the amount of training data remained consistent across all methods.
@@
Finetuning LLMs using PPO is known to be sensitive to hyperparameter selection, and finding the optimal settings is critical for achieving strong performance. To ensure the robustness of our study, we explored hyperparameter values reported in recent studies (Shao et al., 2024; Zheng et al., 2023; Ivison et al., 2024; Huang et al., 2024) and conducted various sweeps across a wide range of values to identify the best configuration for our tasks and models. The full set of hyperparameters, along with their respective search spaces, is detailed in Table 1.
@@
We utilized the same hyperparameter setup as in the PPO implementation (Table 1) for VinePPO. As outlined in Section 5, the number of MC samples, $K$, was set to 9 for all experiments.
@@
C.7 Train vs. Test During Training
When training on reasoning datasets, the training data can be viewed as a finite resource of learning signals. Algorithms that exhaust this resource through memorization tend to generalize less effectively on the test set. As we move from RL-free methods or less accurate credit assignment towards more accurate credit assignment, or full reinforcement learning—from RestEM to DPO, PPO, and finally VinePPO—the model demonstrates higher test accuracy gains per unit of training data consumed. This trend is illustrated in Figure C.4.
Figure C.4: A scatter plot showing the relationship between achieved training accuracy and test accuracy at various checkpoints throughout training. This plot highlights the dynamics of overfitting and generalization across different methods. As we progress from no credit assignment to accurate credit assignment—from RestEM to DPO+, PPO, and finally VinePPO—generalization improves and overfitting decreases. In other words, by treating the training dataset as a resource, VinePPO achieves higher test accuracy per unit of training data consumed. Note that all these are fully trained. Note that the training accuracy does not reach 100 percent due to several factors, including mechanisms like the KL penalty in DPO+, PPO, and VinePPO, the reset to the base model in RestEM, or the absence of any correct self-generated responses for certain questions.
C.8 Compute
@@
C.9 Software Stack
Both PPO and VinePPO require a robust and efficient implementation. For model implementation, we utilize the Huggingface library. Training is carried out using the DeepSpeed distributed training library, which offers efficient multi-GPU support. Specifically, we employ DeepSpeed ZeRO stage 0 (vanilla data parallelism) for RhoMath 1.1B and ZeRO stage 2 (shared optimizer states and gradients across GPUs) for DeepSeekMath 7B . For trajectory sampling during RL training, we rely on the vLLM library (Kwon et al., 2023), which provides optimized inference for LLMs. Additionally, VinePPO leverages vLLM to generate Monte Carlo samples for value estimation. This software stack ensures that our experiments are both efficient and reproducible. For instance, during VinePPO training, we achieve an inference speed of up to 30K tokens per second using 8 $\times$ Nvidia H100 GPUs with the DeepSeekMath 7B model.
C.10 Reproducibility
In this study, all experiments were conducted using open-source libraries, publicly available datasets, and open-weight LLMs. To ensure full reproducibility, we will release both Singularity and Docker containers, equipped with all dependencies and libraries, enabling our experiments to be run on any machine equipped with NVIDIA GPUs, now or in the future. Additionally, we will make our codebase publicly available on GitHub at https://github.com/McGill-NLP/VinePPO.
@@ D.1 Training Plots
D.2 Value Prediction Analysis
In this section, we provide additional plots for value analysis. Specifically, Figures D.8, D.9, D.10 and D.11 demonstrates these plots for on the MATH dataset, and Figures D.12, D.13, D.14 and D.15 on the GSM8K dataset.
Furthermore, we present the prediction error per step in Figures D.16, D.18, D.17 and D.19.
Figure D.5: Comparison of the training behavior between VinePPO and PPO. VinePPO demonstrates consistently higher accuracy throughout the training on the GSM8K dataset. Refer to Figure 5 for MATH dataset.
Figure D.6: Task accuracy as a function of KL divergence during training on the GSM8K dataset. VinePPO significantly higher accuracy per KL. Refer to Figure 6 for MATH dataset.

### v2

@@ 8 Discussion
We showed that better credit assignment improves RL training of LLMs. VinePPO is a stepping stone to identify and fix PPO’s broken credit assignment. It also opens two future research directions. VinePPO is the first RL post-training algorithm that scales generalization slope by scaling post-training compute. Algorithms that have better generalization trends are valuable given the limited resource of truly challenging and verifiable reasoning tasks. Second, VinePPO highlights the value of reconsidering the implicit assumptions behind default algorithm implementations borrowed from Deep RL. In Deep RL, we typically start with a random policy, making it crucial to quickly improve the model’s performance. In this context, it’s more effective to allocate compute toward gathering additional environment samples rather than perfecting each gradient update. However, with an already capable LLM, it is better to spend more compute to make sure we steer its weights carefully. Overall, we hope VinePPO inspires the community to develop more effective RL training algorithms for LLMs.
Impact Statement
Our work aims to improve the ability of large language models to perform complex reasoning tasks, potentially contributing to advances in fields such as education, scientific research, and software development. At the same time, more capable reasoning systems could be used irresponsibly, for instance, by automating sophisticated misinformation or other harmful applications. We therefore encourage researchers and practitioners to employ appropriate safeguards when applying our methods. Overall, this paper advances fundamental techniques in machine learning; its societal impact will depend on responsible deployment and continued ethical considerations by the community.
@@ Appendix B Reasoning Step Separation Examples
Table 2: Summary of RLOO and GRPO hyperparamters used in the experiments.
Parameter Value
Training
Optimizer AdamW
Adam Parameters ($\beta_{1},\beta_{2}$) ($0.9$, $0.999$)
Learning rate $1\times 10^{-6}$
Weight Decay $0.0$
Max Global Gradient Norm for Clipping $1.0$
Learning Rate Scheduler Polynomial
Warm Up $3\%$ of training steps
# Train Steps For MATH dataset $1000$ steps (around 8 dataset epochs)
# Train Steps For GSM8K dataset $650$ steps (around 8 dataset epochs)
General
Maximum Response Length 1024 tokens
Maximum Sequence Length for RhoMath 1.1B 2048 tokens
Maximum Sequence Length for DeepSeekMath 7B 2500 tokens
RL Algorithm
# Responses per Prompt $8$
# Episodes per PPO Step $512$
# Prompts per PPO Step $512/8=64$
Mini-batch Size $64$
# Inner epochs per PPO Step $2$
Sampling Temperature $0.6$
Discount Factor $\gamma$ $1.0$
[…]
Policy Clipping Parameter $\epsilon$ $0.2$
@@ C.4 Evaluation
C.5 Other Baselines
GRPO (Shao et al., 2024) and RLOO (Ahmadian et al., 2024) GRPO replaces PPO’s value network with a policy gradient baseline computed from the average return of a group of responses to the same input. For each training question ${\bm{x}}$, all algorithms generates $G$ responses, yielding training trajectories $\tau_{1},\tau_{2},\ldots,\tau_{G}\sim\pi(\cdot|{\bm{x}})$ with corresponding returns $R_{1},R_{2},\ldots,R_{G}$. Note that in the case of GRPO, we need to have $G>1$. Then, GRPO computes the empirical mean $\mu_{\bm{x}}=\frac{1}{G}\sum_{i=1}^{G}R_{i}$ and standard deviation $\sigma_{\bm{x}}$ of these returns. For each trajectory $\tau_{i}$, the advantage $A(s,a)$ for all state-action pairs $(s,a)\in\tau_{i}$ is defined as:
$A(s,a)=\frac{R_{i}-\mu_{\bm{x}}}{\sigma_{\bm{x}}}.$
Notably, this introduces bias in policy gradient estimation because the return $R_{i}$ of the current trajectory is used in computing its own baseline. RLOO addresses this bias by employing a leave-one-out strategy for baseline computation. Specifically, for each trajectory $\tau_{i}$, the baseline is computed using the returns of all other trajectories in the group, excluding $R_{i}$. Let $\mu_{\bm{x}}^{(i)}$ denote the empirical mean of $\{R_{j}\}_{j\neq i}$. The advantage for all state-action pairs in $\tau_{i}$ is then computed as:
$A(s,a)=R_{i}-\mu_{\bm{x}}^{(i)}.$
This modification ensures that the baseline for each trajectory is independent of its own return, yielding an unbiased policy gradient estimate.
@@ C.5 Other Baselines
Therefore, maximizing log-likelihood training on correct responses is equivalent to train with policy gradient without precise credit assignment, such as without advantages for specific actions. In our experiments, we observe the impact of this limitation in both Figure C.3 and Figure 5 where RestEM overfits on the training data.
@@ C.6 Hyperparameters
In this section, we present a comprehensive overview of the hyperparameters used in our experiments. It’s important to note that the number of training episodes was carefully selected to ensure that the amount of training data remained consistent across all methods.
@@
Finetuning LLMs using PPO is known to be sensitive to hyperparameter selection, and finding the optimal settings is critical for achieving strong performance. To ensure the robustness of our study, we explored hyperparameter values reported in recent studies (Shao et al., 2024; Zheng et al., 2023; Ivison et al., 2024; Huang et al., 2024) and conducted various sweeps across a wide range of values to identify the best configuration for our tasks and models. Specifically, we find the set of hyperparameters that perform best across both MATH and GSM8K using RhoMath 1.1B model. Then, we employ the optimal set of parameters for the rest of our experiments. The full set of hyperparameters, along with their respective search spaces, is detailed in Table 1.
@@
We utilized the same hyperparameter setup as in the PPO implementation (Table 1) for VinePPO.
RLOO and GRPO
Since policy optimization in RLOO and GRPO is similar to PPO, we initialze their hyperparameters from PPO. This not only ensure we start from a strong set of values, but also allows for a systematic comparison among these algorithms. We further tune their KL coefficient for stable training. Note that lack of credit assignment mechanism could lead to high variance policy gradient update, resulting in unstable training (Greensmith et al., 2001). See Table 2 for the full list.
@@
Figure C.4: Comparison of the training behavior between VinePPO and PPO. VinePPO demonstrates consistently higher accuracy (as measured on the test set of MATH dataset) throughout the training. Refer to Appendix D for more detailed plots.
C.7 Compute
@@
C.8 Software Stack
Both PPO and VinePPOrequire a robust and efficient implementation. For model implementation, we utilize the Huggingface library. Training is carried out using the DeepSpeed distributed training library, which offers efficient multi-GPU support. Specifically, we employ DeepSpeed ZeRO stage 0 (vanilla data parallelism) for RhoMath 1.1B and ZeRO stage 2 (shared optimizer states and gradients across GPUs) for DeepSeekMath 7B . For trajectory sampling during RL training, we rely on the vLLM library (Kwon et al., 2023), which provides optimized inference for LLMs. Additionally, VinePPOleverages vLLM to generate Monte Carlo samples for value estimation. Specifically, after each RL training iteration, the current policy’s checkpoint is loaded into vLLM. Then, we use vLLM’s serving API to sample new trajectories and also Monte Carlo Samples for VinePPO’s value estimation. In our setup, we spawn a separate vLLM engine on each GPU rank. This would allow for data parallelism during both sample generation and training. This software stack ensures that our experiments are both efficient and reproducible. For instance, during VinePPO training, we achieve an inference speed of up to 30K tokens per second using 8 $\times$ Nvidia H100 GPUs with the DeepSeekMath 7B model.
C.9 Reproducibility
In this study, all experiments were conducted using open-source libraries, publicly available datasets, and open-weight LLMs. To ensure full reproducibility, we will release both Singularity and Docker containers, equipped with all dependencies and libraries, enabling our experiments to be run on any machine equipped with NVIDIA GPUs, now or in the future. Additionally, we will make our codebase publicly available on GitHub at https://github.com/McGill-NLP/VinePPO
@@ D.1 Training Plots
D.2 Explained Variance and Mean Absolute Error (MAE) of Value Prediction During Training
To ensure healthy training runs, we assess value prediction accuracy using explained variance and mean absolute error (MAE). Explained variance quantifies how much of the variance in ground-truth values is captured by the estimator:
[…]
where $v_{g}$ are ground-truth values, $\hat{v}_{g}$ are predictions, and $\bar{v}=\frac{1}{n}\sum_{g=1}^{n}v_{g}$ is the mean of the ground-truth values. The mean absolute error (MAE) is given by:
$\text{MAE}=\frac{1}{n}\sum_{g=1}^{n}\bigl{|}v_{g}-\hat{v}_{g}\bigr{|}.$
As shown in Figure D.5, PPO shows improving explained variance and decreasing MAE, indicating stable training. VinePPO achieves the highest explained variance and lowest MAE. RLOO and GRPO are included solely for demonstration, illustrating the deviation of their baselines from ground truth value estimates.
Figure D.5: Explained Variance and Mean Absolute Error of values. VinePPO demonstrates higher explained variance in value predictions and lower mean absolute error compared to RLOO, GRPO, and PPO across both datasets. Additionally, PPO’s value predictions show non-negative explained variance values close to one, indicating stable and effective training. Note that RLOO and GRPO are included solely for demonstration, illustrating the deviation of their baselines from ground truth value estimates.
D.3 KL Divergence
The RL objective (Equation 1) balances maximizing task performance while constraining deviations from the initial policy $\pi_{\mathrm{ref}}$, measured by KL divergence. We analyze how VinePPO and PPO navigate this trade-off by plotting task accuracy against KL divergence $\mathrm{KL}[\pi_{\theta}\|\pi_{\mathrm{ref}}]$ throughout training (Figure D.8). Results show VinePPO consistently achieves higher accuracy at same KL divergence, indicating more efficient use of the “KL budget.” This efficiency stems from VinePPO’s more precise credit assignment. As shown in Figure 1, many advantages are zero, and VinePPO excludes these steps from the loss. By avoiding unnecessary updates on non-contributing tokens, VinePPO reduces non-essential parameter adjustments that would inflate KL.
D.4 Temperature Tolerance
Sampling temperature is a critical hyperparameter controlling the randomness of sampled trajectories. At higher temperatures models generates more diverse trajectories, accelerating early training through increased exploration. However, this diversity challenges PPO’s value network, requiring generalization over a wider range of states. We compared VinePPO and PPO using temperatures $T\in\{0.6,0.8,1.0\}$ over the initial third of training steps. Figure D.10 shows VinePPO consistently benefits from higher temperatures, achieving faster convergence. Conversely, PPO fails to benefit from increased exploration and even diverges at $T=1.0$, where trajectories are most diverse.
D.5 Value Prediction Analysis
In this section, we provide additional plots for value analysis. Specifically, Figures D.13, D.14, D.15 and D.16 demonstrates these plots for on the MATH dataset, and Figures D.17, D.18, D.19 and D.20 on the GSM8K dataset.
Furthermore, we present the prediction error per step in Figures D.21, D.23, D.22 and D.24.
Figure D.6: Comparison of the training behavior between VinePPO and PPO. VinePPO demonstrates consistently higher accuracy throughout the training on the GSM8K dataset. Refer to Figure C.4 for MATH dataset.
Figure D.7: Task accuracy as a function of KL divergence during training on the GSM8K dataset. VinePPO significantly higher accuracy per KL. Refer to Figure D.8 for MATH dataset.
Figure D.8: Task accuracy as a function of KL divergence during training on the MATH dataset. VinePPO achieves higher accuracy, reflecting more efficient credit assignment and focused updates.
@@ D.5 Value Prediction Analysis
Figure D.10: Test set accuracy during training with higher temperature presented for DeepSeekMath 7B and MATH dataset. VinePPO can tolerate higher temperatures.
Figure D.11: Ablating the number of auxiliary trajectories $K$ for estimating $\hat{V}_{\mathrm{MC}}(s_{t})$ on RhoMath 1.1B and GSM8K. Increasing $K$ consistently improves task performance. (see Figure 7 for MATH dataset)
Figure D.12: Value prediction accuracy formulated as a classification problem, where a prediction is considered correct if it falls within 0.05 of the ground truth.
@@
Appendix E More Examples of Advantages in VinePPO
In addition to Figure 1, we provide more examples of effective advantage computation of VinePPO in Figures E.25, E.26 and E.27.
Prompt ($s_{0}$) Advantage
The vertex of the parabola described by the equation $y=-3x^{2}-30x-81$ is $(m,n)$. What is $n$?
Response
$s_{1}$$s_{1}$$0.0$ We can re-write the given equation as $y=-3(x^{2}+10x)-81$.
$s_{2}$$s_{2}$$\mathbf{0.2}$ To complete the square, we add $(-10/2)^{2}=25$ inside the parentheses:
$s_{3}$$s_{3}$$\mathbf{0.5}$ $y=-3(x^{2}+10x+25)-81-(-3)(25)$
$s_{4}$$s_{4}$$0.0$ $=-3(x+5)^{2}-81+75$
$s_{5}$$s_{5}$$0.0$ $=-3(x+5)^{2}-6.$
$s_{6}$$s_{6}$$0.0$ The vertex of a parabola in the form $a(x-h)^{2}+k$ is at the point $(h,k)$,
$s_{7}$$s_{7}$$0.0$ so the vertex of the given parabola is at the point $(-5,-6)$.
$s_{8}$$s_{8}$$0.0$ Therefore, $n=\boxed{-6}$.
Example of Advantages Computed by VinePPO
Figure E.25: Steps $s_{2}$ and $s_{3}$ provide a critical insight for advancing the solution since they provide a way to organize the equation.
Prompt ($s_{0}$) Advantage
$0.4$ Let $f(x)=2x-3$ and $g(f(x))=5-4x.$ Find $g(4).$
Response
$s_{1}$$s_{1}$$0.0$ We know that $g(f(x))=5-4x.$
$s_{2}$$s_{2}$$\mathbf{-0.6}$ So, if we can figure out what $f(x)$ is when $x=4,$
$s_{3}$$s_{3}$$0.0$ we can plug that value into $g(f(x))$ to find $g(4).$
$s_{4}$$s_{4}$$0.0$ Since $f(x)=2x-3,$ we have $f(4)=2(4)-3=5.$
$s_{5}$$s_{5}$$0.0$ Therefore, $g(f(4))=g(5)=5-4(5)=5-20=\boxed{-15}.$
Example of Advantages Computed by VinePPO
Figure E.26: This example demonstrates an error in step $s_{2}$ where the correct step would be …what $x$ is when $f(x)=4$. This mistake prevents to model to successfully solve the question after $s_{2}$.
Prompt ($s_{0}$) Advantage
Write the expression $\frac{4+3c}{7}+2$ as a single fraction.
Response
$s_{1}$$s_{1}$$0.0$ We get a common denominator of 7:
$s_{2}$$s_{2}$$0.0$ $\frac{4+3c}{7}+2=\frac{4+3c}{7}+\frac{2\cdot 7}{7}$
$s_{3}$$s_{3}$$\mathbf{-0.3}$ $=\frac{4+3c}{7}+\frac{14}{7}$
$s_{4}$$s_{4}$$\mathbf{0.4}$ $=\frac{4+3c+14}{7}$
$s_{5}$$s_{5}$$0.0$ $=\boxed{\frac{18+3c}{7}}.$
Example of Advantages Computed by VinePPO
Figure E.27: Steps $s_{3}$ and $s_{4}$, despite impacting the probability of solving the question, they are not meaningfully incorrect or insightful.